OpenStudy (anonymous):

P(B|A)=P(B)*P(A|B)/P(A) <---Formula Consider 3 coins, 2 that are fair and 1 that is not. Assume that the unfair coin has a 3/4 probability of landing heads and a 1/4 probability of landing tails. Question: What is the probability that the coin you have chosen out of these 3 is fair, given that you flip 5 heads in a row? I need help understanding this... It is really confusing

OpenStudy (anonymous):

P(B|A) is the probability of B, given A. Put in this context, you might write it this way: P(Coin is fair | flipped 5 heads in a row) = P(Coin is fair) * P(Flipped 5 heads in a row | coin is fair) / P(flipped 5 heads in a row)

OpenStudy (anonymous):

So how would you multiply and devide that? would you use the 3/4 and 1/4 values in the equation?

OpenStudy (anonymous):

In this case, since 1 coin is unfair and 2 are fair, P(Coin is fair) = 2/3 P(flipped 5 heads in a row | coin is fair) = 1/2 * 1/2 * 1/2 * 1/2 * 1/2 That's the easiest part. The only mildly confusing part would be P(flipped 5 heads in a row) but we can find that this way: P(flipped 5 heads in a row) = 2/3 * ( 1/2 * 1/2 * 1/2 * 1/2 *1/2) + 1/3 * (3/4 * 3/4 * 3/4 * 3/4 * 3/4)

OpenStudy (anonymous):

In the last part, the probability is like a weighted average -- 2/3 from the fair coins, and 1/3 from the unfair coin.

OpenStudy (anonymous):

okay I see

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Sure, no problem.